# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

### v1

@@ Abstract
Deep learning for human sensing on edge systems offers significant opportunities for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. Current methods that rely on transferring pre-trained models often encounter issues such as modality shift and high resource demands, resulting in substantial accuracy loss, resource overhead, and poor adaptability across different sensing applications. In this paper, we propose XTransfer, a first-of-its-kind method for resource-efficient, modality-agnostic model transfer. XTransfer freely leverages single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely repairs modality shift in pre-trained model layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance on human sensing tasks while significantly reducing the costs of sensor data collection, model training, and edge deployment.
@@ 1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (humansensing2023, ; Ystgaard2023, ). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed2018, ). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Xiaoqiang2019, ; zhang2020, ; Yanjiao2020, ). However, existing DL solutions are often data- and resource-intensive for training and deployment (ondevicedl, ), posing significant challenges for data collection and edge deployment in human sensing (surveyiot, ; iotdatachallenge, ).
@@
Unlike other data modalities (e.g., text or image), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane2015, ; Jeyakumar19, ; Teh2020, ). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Kroger2019, ; zhangipsn, ). Manual annotation of large volumes of sensor data introduces additional complexity and incurs high costs (Song2023fewshot, ; chen2020deep, ). Furthermore, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), thereby adding complexity to model training, adaptation, and edge deployment (humansurvey, ; humansensing2023, ).
Prior research has explored Few-shot Learning (FSL) (metasense, ; Wang2024, ), transfer learning (tlhar, ), and their combination (i.e., cross-domain FSL) (crossHAR, ; FewSense, ) to enable DL-based human sensing with few data. However, they either still require large-scale labeled datasets from the same singular modality to costly train source models from scratch (Song2023fewshot, ), or face challenges such as modality shift and limited adaptability when leveraging pre-trained models (Jaehoon2022, ; Ran2022, ; Sommelier2022, ), particularly in cross-modality settings.
@@
These limitations highlight the need for a scalable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data. Recent advances in multi-modality learning (mmfi, ; chen2025xfi, ; chen2024probabilistic, ; madaan2024jointly, ; yang2024facilitating, ; baldenweg2024multimodal, ) with methods such as knowledge distillation (e.g., image-to-sensor transfer (zhao2018through, ; Xingzhe2022, ; Sevgi2020, ; Yu2023, )) and multi-modal model (e.g., CLIP-based models (girdhar2023imagebind, ; wu2023nextgpt, )), demonstrate the growing feasibility of cross-modality model transfer, driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (PyTorchhub, )), which opens new opportunities for real-world human sensing. However, due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Sommelier2022, ; Zhang2023, ; Jaehoon2022, ). Our experiments reveal that state-of-the-art baselines (Tab. 1), though effective on vision and NLP tasks, fail to reach oracle performance (metasense, ) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Fig. 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Sicheng2020, ; Xiangyumfs, ; NEURIPS20196048ff4e, ). While methods such as neural architecture search (NAS) (searchinsights, ; AdaptiveNet2023, ; LegoDNN2021, ), pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), or quantization (Polino2018ModelCV, ), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Fig. 2(c), our experiments using structural pruning (frankle2018the, ; NEURIPS2022, ) demonstrate significant model accuracy loss (50.5% on average). These challenges point to a critical gap in developing adaptable and scalable methods.
@@
To overcome the challenges, we propose XTransfer, a first-of-its-kind cross(X)-modality model Transfer method that freely leverages public source models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks, significantly reducing the need for extensive sensor data collection and training from scratch. It also offers resource-efficient designs to optimize edge deployment, paving the way for scalable real-world applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained source models under cross-modality FSL settings (see App. A), resulting in compact and adaptable models. To mitigate modality shift and prevent overfitting, it enables model repairing achieved by a novel Splice-Repair-Removal (SRR) pipeline (c.f. Sec. 4). This pipeline transforms a few sensor data to broadly adapt pre-trained model layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (c.f. Sec. 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, this mechanism discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (searchinsights, ). To further accelerate the search when dealing with multiple sources, it integrates a pre-search check strategy to efficiently avoid unnecessary repairing.
Comprehensive evaluations (c.f. Sec. 6) are conducted on 8 public datasets along with their corresponding pre-trained source models, spanning diverse modalities (e.g., image, text, audio, and sensing), 4 collected target sensing datasets in real-world settings, 2 public sensing datasets and 1 standard image dataset as targets (Tab. 2) on 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (c.f. Sec. 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while using few data with broad adaptability. XTransfer also streamlines the layer search process, reducing search time by 2.1 to 4 times. The main contributions of this work are:
  • A first-of-its-kind method, XTransfer, that enables resource-efficient, modality-agnostic model transfer for human sensing with few data at the edge, as shown in Fig. 3.
  • A novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities.
  • A novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner, creating compact models.
  • Extensive experimental evaluations, benchmarking state-of-the-art baselines based on various source and target datasets, validating competitive scalable performance of XTransfer.
@@
Figure 1: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet and HHAR dataset as target under a 5-shot setting (see App. A).(overfitting, ). (b) shows the average similarity and FSL difficulty across all sensing datasets (Tab. 2) to each source modality (e.g., Image, Text, Sensing) (c.f. Sec. 3). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Sec. 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability (c.f. Sec. 4). (c) shows model optimisation bottlenecks.
@@ 2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based loss functions (Zhao2021, ; Yue2021Prototypical, ), whereas multi-source methods leverage the Wisdom of the Crowd principle through source distillation (Sicheng2020, ; Xiangyumfs, ) and maximal correlation analysis (NEURIPS20196048ff4e, ). Existing attempts on human sensing tasks (metasense, ; Wang2024, ; tlhar, ; crossHAR, ; FewSense, ) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and model training with limited adaptability. Recent efforts have also explored fine-tuning large language models (e.g., GPT2 (zhou2023one, )) for time-series analysis. However, testing on benchmarks (Boosting2021, ; Jaehoon2022, ; chen2018a, ), our results reveal that they still suffer from severe overfitting under modality shift and few data. Also, none of them explicitly considers resource constraints of edge systems. In contrast, XTransfer enables model repairing to safely mitigate modality shifts using only few sensor data, enabling modality-agnostic model transfer with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), quantization using low-precision data types (Polino2018ModelCV, ), model merging by fine-tuning shared layers (Gemel2023, ), one-size-fits-all model (cai2020once, ), and NAS techniques such as model structure optimization (AdaptiveNet2023, ; LegoDNN2021, ) or search using few data (neuralFT, ; xu2022fewshot, ), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures by efficiently searching and recombining layers of interest from pre-trained source models. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
@@ 3 Preliminary
Insights in cross-modality. We apply a standard cross-domain FSL benchmarking (Jaehoon2022, ) to gain insights. It uses the Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets, including 5 human sensing target datasets and 2 well-known image datasets (Tab. 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Fig. 2(b).
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (islam2021broad, ), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (i.e., channel importance) (pmlrv162luo22c, ) of each layer to indicate layer-wise performance. We define a single layer or a dependent layer block (i.e., L-Blocks in Tab. 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Fig. 2(a) shows that negative transfer (Zhang2023, ) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Fig. 2(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (shahapure2020cluster, ), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts, but several technical challenges arise. First, the pre-trained feature space at each layer is highly complex due to the multidimensional nature of MMC, and this complexity increases with layer depth. Such complexity poses challenges for "repairing" latent feature distributions across modalities using few sensor data. Second, strong layer-wise dependencies, formed during source model training, become increasingly rigid in deeper layers, posing challenges for standard techniques (e.g., N-Pair loss (Sohn2016, ), Triplet loss (Schroff2015, )) that may fail to preserve these dependencies. Third, Fig. 2(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence, indicating that not all repaired layers contribute positively. This motivates our LWS control design, which selectively recombines effective layers (i.e., layers of interest) while discarding others. However, such selection must account for on-device resource constraints to optimize the layer growth. The search space also expands rapidly with the number of layers or source models, further complicating efficient search.

### v2

# XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge

@@ Abstract
Deep learning for human sensing on edge systems presents significant potential for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and poor adaptability in practice. In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design. XTransfer flexibly uses single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
@@ 1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a; Ystgaard et al., 2023). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed & Farid, 2018). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020). However, existing DL solutions are often data- and resource-intensive for training and deployment (Dhar et al., 2021), posing significant challenges for data collection and edge deployment in human sensing (Sobin, 2020; Liu et al., 2019).
@@
Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane & Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022). Manual annotation of large volumes of sensor data is labor-intensive and incurs high costs (Song et al., 2023). Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
Prior research has explored Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane & Ploetz, 2024), and their combination (i.e., cross-domain FSL) (Thukral et al., 2025; Yin et al., 2024) to notably reduce data collection costs by adapting pre-trained models to sensing applications. However, they still rely on large-scale labeled datasets from the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift and poor adaptability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings. Recent advances in multi-modal learning (Yang et al., 2023; Chen & Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), demonstrate the feasibility of knowledge transfer across modalities, but they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge. These limitations motivate the need for a scalable and adaptable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
@@
In practice, however, addressing this need is uniquely challenging. Due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022). Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019). While methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle & Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
Figure 1: Overview. XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control). LWS control first segments source models into layers and operates layer-wise search across pools. At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest. These layers are incrementally recombined during the search, resulting in a compact model for enabling human sensing at the edge. Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
@@
Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer. It flexibly leverages public pre-trained models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few sensor data, significantly reducing the costs of large-scale data collection and training from scratch. Its resource-efficient design further optimizes edge deployment, enabling both scalability and adaptability to diverse human sensing applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained models under cross-modality FSL settings (see Appendix A), resulting in compact models. To mitigate modality shift and prevent overfitting, it enables model repairing using a novel Splice-Repair-Removal (SRR) pipeline (see Section 4). SRR pipeline transforms a few sensor data to broadly adapt pre-trained layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (see Section 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, LWS control discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (White et al., 2023). To further accelerate the search in multi-source settings, a pre-search check strategy is designed to efficiently avoid unnecessary repairing.
The main contributions are summarized as: 1) a first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems, as shown in Figure 1; 2) a novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities; 3) a novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner to create compact models; and 4) extensive experimental evaluations (see Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on 8 public datasets with corresponding pre-trained source models across diverse modalities (e.g., image, text, audio, and sensing) and 7 target datasets (Table 2) using 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (see Section 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while reducing layer search time by 2.1 to 4 times and substantially lowering the costs of data collection, model training, and edge deployment, demonstrating broad adaptability.
@@
Figure 2: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting (see Appendix A).. (b) shows the average similarity and FSL difficulty across all sensing datasets to source modalities (e.g., Image, Text, Sensing) using default reshaping (see Appendix B.1). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Section 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability. (c) shows model pruning bottlenecks.
@@ 2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based objectives (Zhao et al., 2021; Yue et al., 2021b), whereas multi-source methods leverage the Wisdom of the Crowd principle through unsupervised source distillation (Zhao et al., 2020; Yue et al., 2021a) or maximal correlation analysis (Lee et al., 2019). Recent efforts have also explored fine-tuning LLMs (e.g., GPT2 (Zhou et al., 2023)) for time-series analysis. However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings. LLM-based data augmentation (Leng et al., 2024; 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus. Existing attempts on human sensing tasks (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training with limited adaptability. In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Frankle & Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model structure optimization (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures while strengthening layer-wise dependence, by efficiently searching and recombining layers of interest from pre-trained source models at scale. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
@@ 3 Preliminary
Insights in cross-modality FSL. We apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights. It uses Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets including 5 human sensing and 2 well-known image datasets (Table 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Figure 2(b).
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (Islam et al., 2021), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (Luo et al., 2022) of each layer as the default metric to quantify layer-wise performance, as it offers a lightweight measure for channel importance. We define a single layer or a dependent layer block (i.e., L-Blocks in Table 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Figure 3(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Figure 3(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (Shahapure & Nicholas, 2020), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts under cross-modality FSL settings, but several technical challenges arise. First, it is hindered by structural and semantic gaps between sources and target data, where classes and semantics often differ. Second, the pre-trained feature space at each layer is highly complex due to the multidimensional and noisy nature of MMC across source models from different modalities. This complexity increases with layer depth, as the strong layer-wise dependence formed during source model training becomes increasingly rigid, making it difficult to repair layer-wise misalignment using only few sensor data. Third, Figure 3(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence. It indicates that not all repaired layers contribute positively, motivating our LWS control design to selectively recombine layers of interest while discarding others. The selection must also account for on-device resource constraints to control layer growth, while the rapidly expanding search space with increasing layers and source models further challenges the layer search.